OpenStudy (winsteria):

please help medal and fan! What is the solution to the equation 2(m-8)=4(m+6) ? m = ?

OpenStudy (winsteria):

@Nnesha @TheSmartOne @KamiBug

TheSmartOne (thesmartone):

Distributive property

Nnesha (nnesha):

first distribute bracket by 2

Nnesha (nnesha):

:P

TheSmartOne (thesmartone):

\(\Huge a(b-c)=ab-ac\)

Nnesha (nnesha):

what is 2 times m = ??? 2 times -8 = ??

TheSmartOne (thesmartone):

and \(\Huge a(b+c)=ab+ac\)

TheSmartOne (thesmartone):

\(\Huge 2(m-8)=?\)

OpenStudy (winsteria):

2 times -8 = -16

Nnesha (nnesha):

what about 2 times m ???

TheSmartOne (thesmartone):

^^

OpenStudy (winsteria):

I don't know what m is...

TheSmartOne (thesmartone):

So for example 5 times x would just be 5x

OpenStudy (winsteria):

Oh um

OpenStudy (winsteria):

2m? I'm confused idk

TheSmartOne (thesmartone):

yes it is 2m!!!

Nnesha (nnesha):

yes

TheSmartOne (thesmartone):

So now we have \(\Huge 2m-16=4(m+6)\)

Nnesha (nnesha):

 now do same thing on right side